OpenStudy (anonymous):

Which value is a solution to the open sentence? 6x – 3 = 39 A. 4 B. 6 C. 5 D. 7

OpenStudy (arthur_ser):

6x-3=39 6x=42 x=42/6 x=7

OpenStudy (anonymous):

can you help me with a few more?